# Jets-separation thresholds, Seshadri constants and higher Gauss-Wahl maps on abelian varieties Nelson Alvarado N. Alvarado

### v1

@@ Abstract.
Given a closed subscheme of a polarized abelian variety we define its vanishing threshold with respect to and relate it to the Seshadri constant of the ideal defining As a particular case, we introduce the notion of jets-separation thresholds, which naturally arise as the vanishing threshold of the -infinitesimal neighborhood of a point. Afterwards, by means of Fourier-Mukai methods we relate the jets-separation thresholds with the surjectivity of certain higher Gauss-Wahl maps. As a consequence we obtain a criterion for the surjectivity of those maps in terms of the Seshadri constant of the polarization
Partially supported by the MIUR Excellence Department Project MatMod@TOV awarded to the Department of Mathematics of the University of Rome Tor Vergata.
1. Introduction
In this article we introduce the vanishing threshold of a closed subscheme of a polarized abelian variety defined over an algebraically closed field of characteristic zero. In the case of a closed point (with its reduced structure) this number is no other than the basepoint-freeness threshold introduced by Jiang-Pareschi in [cohrank], which has also been studied by Caucci [caucci], Ito [itobir] and Jiang [jiang], using methods from birational geometry, and by Rojas [Rojas], by means of stability conditions in the two-dimensional case. When we consider the -infinitesimal neighborhood of a closed point (i.e the scheme defined by the -power of the ideal of the point) a natural notion of -jets separation threshold arise.
@@ 1. Introduction
Subsequently, we show that -jets separation thresholds are intimately related, via the Fourier-Mukai transform, to certain higher Gauss-Wahl maps. In fact, we prove that they carry esentially the same information of certain (suitably defined) -Gauss-Wahl surjectivity threshold. This paralells the fact, shown by Pareschi-Jiang in [cohrank, Proposition 8.1], that basepoint freeness threshold carries the same information as the surjectivity threshold for multiplication maps of global sections (and therefore, is intimately related to projective normality). As a byproduct of the aforementioned relation we obtain a description of the Seshadri constant in terms of the asymptotic behaviour of the surjectivity of certain Gauss-Wahl maps. Finally, by means of Nadel’s vanishing theorem, we bound the -jets separation thresholds in terms of the Seshadri constant, when we work over the complex numbers. As a consequence, we establish precise lower bounds for the Seshadri constant guaranteeing the surjectivity of higher Gauss-Wahl maps, asymptotically improving the results present in the literature.
@@
Now, let be a polarization on (that is, an ample class in ). As shown by Pareschi-Jiang, the IT(0) condition can be extended for -twisted sheaves (we recall this notion in Section 2.1). In this context, we define the vanishing threshold of with respect to as the real number
@@
Our first result relates the vanishing threshold of infinitesimal neighborhoods of a closed subscheme with the Seshadri constant of the ideal defining More precisely, we prove the following:
@@
For a complex abelian variety the following inequalities of jets separation thresholds hold:
@@
For instance, the above result says that if that is, if separates -jets at every point, then is surjective, while is already surjective whenever On the other hand, using Theorem 1.1 3) we obtain a condition ensuring the surjectivity of certain Gauss-Wahl maps in terms of the Seshadri constant, which (at least asymptotically) improves the results present in the literature ([mult, Theorem 2.2], for instance). More precisely, we get the following:
@@
This article structures as follows: in Section 2 we survey the notions of generic vanishing for -twisted objects and their relation with the Fourier-Mukai functor, whose definition we also recall. In Section 3 we formally introduce the notion of vanishing threshold of a closed subscheme of an abelian variety and, in particular, the -jets separation thresholds. In this context, we establish Theorem 1.1 2.a). In Section 4 we discuss Seshadri constants of ideal sheaves, and relate them with the vanishing threshold of the corresponding closed subscheme, proving Theorem 1.11) and 2.b). In Section 5 we discuss Gauss-Wahl maps, introduce their surjectivity thresholds and prove a slightly more general version of Theorem 1.2 above by computing the Fourier transform of In Section 6, we prove Theorem 1.1 3) and show its consequences. In particular, in this section we prove Theorem 1.4 and Corollary 1.5. Finally, in Section 7 we state a couple of questions for future research: the first one regards the possibility to render our asymptotic results more effective and the second one regards the possibility to extend Theorem 1.1 3) to abelian varieties defined over fields of positive characteristic.
Acknowledgements: This work is part of my Phd thesis. I would like to thank my advisor, Giuseppe Pareschi, for his invaluable guidance and help to improve both the exposition and the mathematical content of this article.
@@ 2. Preliminaries
In this section we recall the preliminar material that we will need. We start by fixing some notation. Let be an abelian variety defined over an algebraically closed field. By a polarization on we mean an ample class For as before, we write for the dual of that is, is the abelian variety parametrizing traslation-invariant line bundles on For a polarization we write (or ) for the isogeny which sends a point to the point of corresponding to the line bundle where is the traslation map given by and is a (any) line bundle representing . For a positive integer we write for the isogeny Finally, we fix a normalized Poincaré bundle on and for write for the line bundle on parametrized by (equivalently, is the fiber of over ).
@@
For a -twisted object it is possible to define its cohomological support loci as follows. Let be a line bundle representing and consider with and Write
where stands for the sheaf (hyper)cohomology. Now, if is another representant of then is just a traslation of On the other hand we note that if we change for with a positive integer and we consider a symmetric , then:
where, to get the direct sum we used [Mumford, p.72] and the projection formula.
In this context, the next definition naturally arise:
@@
For a -twisted object its -th cohomological support loci is the class of where is a representant of and under the equivalence relation generated by translations and by taking direct images of multiplication isogenies.
In particular, although is not a set, its dimension and codimension are well defined and are denoted by and respectively. In [cohrank, Section 5] the cited authors use these loci in order to define generic vanishing notions for -twisted objects:
@@
For example, if is a polarization, then from Kodaira vanishing it follows that is IT(0). Another important example of these generic vanishing conditions regards the notion of separation of jets that we now recall:
@@
In particular, separate -jets at if and only if is not a basepoint of Furthermore, if is not a basepoint of then separates 1-jets at if and only if separates tangent vectors at (i.e the map between tangent spaces induced by the rational morphism is injective).
@@
It follows that is GV if and only if there exist such that separate -jets at and it is IT(0) if and only if this happens for all In particular, is always GV while it is IT(0) if and only if is basepoint free.
We have the following characterization of the behaviour of the generic vanishing conditions of for varying
@@ 2.2. Fourier-Mukai transform
In particular, we see that if is IT(0) then for every
@@
Moreover, we have the following equivalence [PPo3, Lemma 3.6] we have that -twisted object is GV if and only if
for a (any) representation and for a (any) representant of
@@ 2.3. Cohomological rank functions
For varying the fact that whether or not, can be expressed using the cohomological rank functions introduced by Pareschi-Jiang in [cohrank]. In this subsection we recall their definition and fix some notation
@@ 3. Vanishing thresholds
In this section we introduce the notion of vanishing thresholds for coherent sheaves. Afterwards, we give a some examples and at the end we prove Theorem 1.1 1) from the introduction.
@@
Let be a coherent sheaf on an abelian variety and be a polarization on The vanishing threshold of with respect to is the real number
@@
We note that, by definition, we have that
@@
We have that is the basepoint-freeness threshold considered in [cohrank, Section 8]. In particular and if and only if has base points. More generally, from Proposition 2.6 we see that
if and only if there exist a representant of such that separates -jets at 0. Arguing as in Example 2.5, we see that this is equivalent to say that for any representant of there exist a point such that separates -jets at such a point.
@@
Consider a principal polarization and a symmetric divisor representing it. If is indecomposable, then fails to separate 1-jets (i.e tangent vectors) just in the 2-torsion points of That is, is GV but not IT(0) which, by Remark 2.10 and Proposition 2.6, means that From [cav, p.99], the same conclusion holds if is decomposable.
@@
Let be a symmetric divisor representing We can see that if we identify with via then for while (for this example the value of does not matter). In this case, the transformation formula (Proposition 2.11b)) and the degeneration of the spectral sequence computing the hypercohomology says that
and thus
@@
It follows that, for any fixed there exist such that for every the restriction map
is not surjective and thus can not be GV. In other words, for any there exist such that and hence the sequence is unbounded.
@@
and spectral sequence already degenerates at the second page. It follows that is a subquotient of and thus
@@
As the at the right is zero for and hence we get the contention
@@
Let such that both and are GV. By 2.7 we have that is GV. If we write and the latter condition means that is GV. Now, as pullbacks commutes with derived tensor products and is a flat morphism (so and are also ideals), the Lemma implies that is GV, as we wanted to see.
Finally, from Feteke’s subadditivity lemma (see, [mustata, Lemma 1.4], for example), it follows that the sequence converges and
@@
Note that using the above proposition we see that if then and thus separates -jets at every point. In particular we recover the well known fact ([Bauer, Theorem 1],[RegAVIII, Theorem (4)]) that for every we have that separates -jets at every point.
@@ 4. Seshadri constants
Let be an ample line bundle on an abelian variety and the corresponding polarization. Let be a closed subscheme of Then:
@@
The proof of this statement closely follows [ELC, Theorem 3.2]. First we prove that In order to do this we consider a rational number and we need to show that there exist (possibly very big) such that for which is enough to find such that
To do this, we start by noticing that by [ELC, Lemma 3.3] there exist such that
for all all and all line bundles where and is the corresponding projection. Now, let such that For all such consider and thus we have
so, in order to get the vanishing (4.1), we need to ensure the vanishing of each direct summand.
@@
Now, as we have that is ample and thus, as we just need to consider finite ’s, by Serre’s vanishing we can take such that the desired vanishings (4.1) hold. Summarizing, we have established that is GV for for and thus
@@
where being -regular with respect to means that for . Now, by definition, we have that
and hence
@@
and therefore, passing to the limit we get
@@ 5. Gauss-Wahl maps
In particular, the right arrow is surjective if and only if generates -jets at every point of In this way we also see that and is the kernel of the evaluation map (which in [cohrank] is denoted by ).
@@
where the surjectivity at the bottom follows from the fact that the sheaf is supported on the diagonal (for example, see [Ricolfi, Proposition 1.3] for details). In particular, from the 5-lemma it follows that is surjective as soon as the evaluation map is surjective, that is, as soon as generates -jets at every point of
@@
Let be an ample line bundle on an abelian variety of dimension Then:
@@
for If separate -jets at every point then for
c)
If separates -jets at every point then we have for every and
@@
a) Write with Without loss of generality we may assume that is symmetric and hence Applying to the exact sequences
@@
b) Write with and As (because ), from (5.1) we get an inclusion
with equality whenever separates -jets at every point. In particular, to prove the first part of b) it is enough to show that the group at the right vanishes when Now, again by (5.4), inductively we get that
as we wanted to see. In the case that separates -jets, in (5.6) there is equality and thus the result follows from (5.5)
c) Now, write with positive integers. As separates -jets at every point we get that
The result then follows from part a).
@@
Note that, when separates -jets at every point, the above lemma tell us that is the GV-threshold (see Definition 3.1). Note also that the invariant is already considered in [cohrank], where it is denoted by In the cited reference it is shown that there is a relation between such number and the base-point freeness threshold, when is globally generated. In the following we compute to show that there is an analagous relation between and
@@
and if and are numerically equivalent then
@@
By the transormation formula (Proposition 2.11) it follows that for we have
@@
From Remark 3.10 we see that if is an ample line bundle then generates -jets at every point. Therefore, by Theorem 5.5 we have
@@
In particular the limit exist as an element of and by Proposition 4.1 it follows that
@@
There exist a principally polarized elliptic curve and a -dimensional polarized abelian variety such that
@@
The equivalence between a) and b) follows as a combination of Proposition 3.7 and Theorem 4.1. From Corollary 5.8 above, the equivalence a)c) is immediate. Finally, the equivalence a)d) is the content of Nakamaye’s theorem [Nakamaye, Theorem 1.1].

### v2

@@ Abstract.
Given a closed subscheme of a polarized abelian variety we define its vanishing threshold with respect to and relate it to the Seshadri constant of the ideal defining As a particular case, we introduce the notion of jets-separation threshold, which naturally arises as the vanishing threshold of the -infinitesimal neighborhood of a point. Afterwards, by means of Fourier-Mukai methods we relate the jets-separation thresholds with the surjectivity of certain higher Gauss-Wahl maps. As a consequence we obtain a criterion for the surjectivity of those maps in terms of the Seshadri constant of the polarization
In this article we introduce the vanishing threshold of a closed subscheme of a polarized abelian variety defined over an algebraically closed field. In the case of a closed point (with its reduced structure) this number is no other than the basepoint-freeness threshold introduced by Jiang-Pareschi in [cohrank], which has also been studied by Caucci [caucci], Ito [itobir] and Jiang [jiang], using methods from birational geometry, and by Rojas [Rojas], by means of stability conditions in the two-dimensional case. When we consider the -infinitesimal neighborhood of a closed point (i.e the scheme defined by the -power of the ideal of the point) a natural notion of -jets separation threshold arises.
@@
Subsequently, we show that -jets separation thresholds are intimately related, via the Fourier-Mukai transform, to certain higher Gauss-Wahl maps. In fact, we prove that they carry esentially the same information of certain (suitably defined) -Gauss-Wahl surjectivity threshold. This paralells the fact, shown by Jiang-Pareschi in [cohrank, Proposition 8.1], that basepoint freeness threshold carries the same information as the surjectivity threshold for multiplication maps of global sections (and therefore, is intimately related to projective normality). As a byproduct of the aforementioned relation we obtain a description of the Seshadri constant in terms of the asymptotic behaviour of the surjectivity of certain Gauss-Wahl maps. Finally, by means of Nadel’s vanishing theorem, we bound the -jets separation thresholds in terms of the Seshadri constant, when we work in characteristic zero. As a consequence, we establish precise lower bounds for the Seshadri constant guaranteeing the surjectivity of higher Gauss-Wahl maps, asymptotically improving the results present in the literature.
@@
Now, let be a polarization on (that is, an ample class in ). As shown by Jiang-Pareschi, the IT(0) condition can be extended for -twisted sheaves (we recall this notion in Section 2.1). In this context, we define the vanishing threshold of with respect to as the real number
@@
Our first result relates the vanishing threshold of infinitesimal neighborhoods of a closed subscheme with the Seshadri constant of the ideal defining (see section 4 for its definition). More precisely, we prove the following:
@@
If the base field has characteristic zero, the following inequalities of jets separation thresholds hold:
where is the Seshadri constant of
@@
For instance, the above result says that if that is, if separates -jets at every point, then is surjective, while is already surjective whenever On the other hand, using Theorem 1.1 3) we obtain a condition ensuring the surjectivity of certain Gauss-Wahl maps in terms of the Seshadri constant, which (at least asymptotically) improves the results present in the literature ([mult, Theorem 2.2], for example). More precisely, we get the following:
@@
This article structures as follows: in Section 2 we survey the notions of generic vanishing for -twisted objects and their relation with the Fourier-Mukai functor, whose definition we also recall. In Section 3 we formally introduce the notion of vanishing threshold of a closed subscheme of an abelian variety and, in particular, the -jets separation thresholds. In this context, we establish Theorem 1.1 2.a). In Section 4 we discuss Seshadri constants of ideal sheaves, and relate them with the vanishing threshold of the corresponding closed subscheme, proving Theorem 1.1.1) and 2.b). In Section 5 we discuss Gauss-Wahl maps, introduce their surjectivity thresholds and prove a slightly more general version of Theorem 1.2 above by computing the Fourier transform of In Section 6, we prove Theorem 1.1 3) and show its consequences. In particular, in this section we prove Theorem 1.4 and Corollary 1.5. Finally, in Section 7 we state a couple of questions for future research: the first one regards the possibility to render our asymptotic results more effective and the second one regards the possibility to extend Theorem 1.1 3) to abelian varieties defined over fields of positive characteristic.
@@ 2. Preliminaries
In this section we recall the preliminary materials that we will need. We start by fixing some notation. Let be an abelian variety defined over an algebraically closed field . By a polarization on we mean an ample class For as before, we write for the dual of that is, is the abelian variety parametrizing traslation-invariant line bundles on For a polarization we write (or ) for the isogeny which sends a point to the point of corresponding to the line bundle where is the traslation map given by and is a (any) line bundle representing . For a positive integer we write for the isogeny Finally, we fix a normalized Poincaré bundle on and for write for the line bundle on parametrized by (equivalently, is the fiber of over ).
@@
For a -twisted object we would like to define its cohomological support loci To start, let be a line bundle representing and consider with and Write
where stands for the sheaf (hyper)cohomology. These subsets depend on the choices of and For instance, if is another representant of then is just a traslation of On the other hand, if we change for with a positive integer not divisible by and we consider a symmetric , then:
where, to get the direct sum we used [Mumford, p.72] and the projection formula. However, we can make the following:
@@
For a -twisted object its -th cohomological support loci is the class of where is a representant of and are integers with and , under the equivalence relation generated by translations and by taking direct images of prime to the characteristic multiplication isogenies (however, see Section 2.2 below).
In particular, although is not a set, its dimension and codimension are well defined and they are denoted by and respectively. In [cohrank, Section 5] the cited authors use these loci in order to define generic vanishing notions for -twisted objects:
@@
For example, if is a polarization, then from the index theorem (see [Mumford, III.16], or Kodaira vanishing, in the caracteristic zero case) it follows that is IT(0) for . Another important example of these generic vanishing conditions regards the notion of separation of jets that we now recall:
@@
In particular, separates -jets at if and only if is not a basepoint of Furthermore, if is not a basepoint of then separates 1-jets at if and only if separates tangent vectors at (i.e the map between tangent spaces induced by the rational morphism is injective).
@@
It follows that is GV if and only if there exists such that separates -jets at and it is IT(0) if and only if this happens for all In particular, is always GV while it is IT(0) if and only if is basepoint free.
We have the following characterization of the behaviour of the generic vanishing conditions of for varying :
@@ 2.2. Fourier-Mukai transform
In particular, we see that if is IT(0) then for every Moreover, we have the following equivalence ([PPo3, Lemma 3.6]):
On the other hand, if is a positive integer (not necessarily coprime to the characteristic of the base field), then by [Mukai, (3.4)] we have that
In particular, as is finite, we see that the following conditions are equivalent:
a)
is GV
b)
for every and a (any) representation
c)
for every and a (any) representation
@@ 2.3. Cohomological rank functions
For varying the fact that whether or not, can be expressed using the cohomological rank functions introduced by Jiang-Pareschi in [cohrank]. In this subsection we recall their definition and fix some notation
@@
Moreover, in Theorem 3.2 from loc-cit it is also shown that cohomological rank functions can be extended to continous functions
@@ 3. Vanishing thresholds
In this section we introduce the notion of vanishing thresholds for coherent sheaves. Afterwards, we give some examples and at the end we prove Theorem 1.1 1) from the introduction.
@@
Let be an abelian variety and Let be a polarization on The vanishing threshold of with respect to is the real number
@@
We note that, by definition, we have that Also, thanks to Theorem 2.6 we have the following equality:
@@
We have that is the basepoint-freeness threshold considered in [cohrank, Section 8]. In particular and if and only if has base points. More generally, from Theorem 2.6 we see that
if and only if there exists a representant of such that separates -jets at 0. Arguing as in Example 2.5, we see that this is equivalent to say that for any representant of there exists a point such that separates -jets at such a point.
@@
Consider a principal polarization and a symmetric divisor representing it. If is indecomposable, then fails to separate 1-jets (i.e tangent vectors) just in the 2-torsion points of That is, is GV but not IT(0) which, by Remark 2.10 and Theorem 2.6, means that From [cav, p.99], the same conclusion holds if is decomposable.
@@
Let be a symmetric divisor representing and identify with via We claim that for while To see this, notice that and hence we have an exact sequence
As and we obtain that for and an exact sequence
It follows that Now, for this example, the value of does not matter (however, see [GV-min-GV, Corollary 4.3]). Let By the transformation formula (Proposition 2.11b)) and the degeneration of the spectral sequence computing the hypercohomology we have that
From our previous calculation we obtain then that
As clearly for we conclude that
@@
It follows that, for any fixed there exists such that for every the restriction map
is not surjective and thus can not be GV. In other words, for any there exists such that and hence the sequence is unbounded.
@@
and the spectral sequence already degenerates at the second page. It follows that is a subquotient of and thus
@@
As the at the right is zero for and hence we get the inclusion
@@
Let such that both and are GV. By 2.7 we have that is GV. If we write and the latter condition means that is GV. Now, as pullbacks commutes with derived tensor products and is a flat morphism (so and are also ideals), Lemma 3.8 above implies then that is GV, as we wanted to see.
Finally, from Fekete’s subadditivity lemma (see, [mustata, Lemma 1.4], for example), it follows that the sequence converges and
@@
Note that using the above corollary we see that if then and thus separates -jets at every point. In particular we recover the well known fact ([Bauer, Theorem 1],[RegAVIII, Theorem (4)]) that for every we have that separates -jets at every point.
@@ 4. Seshadri constants
Let be an ample line bundle on an a -abelian variety and the corresponding polarization. Let be a closed subscheme of Then:
@@
The proof of this statement closely follows [ELC, Theorem 3.2]. First we prove that In order to do this we consider a positive rational number with (note that this kind of numbers form a dense subset of ) and we need to show that there exists (possibly very big) such that for which is enough to find such that
To do this, we start by noticing that by [ELC, Lemma 3.3] there exists such that
for all all and all line bundles where and is the corresponding projection. Now, let be an integer such that For all such consider and thus we have
where in the last step we used the fact that to get the direct sum ([Mumford, p.72]). So, in order to get the vanishing (4.1), we need to ensure the vanishing of each direct summand.
@@
Now, as we have that is ample and thus, as we just need to consider finite ’s, by Serre’s vanishing we can take such that the desired vanishings (4.1) hold. Summarizing, we have established that is IT(0) for and thus
@@
where being -regular with respect to means that for . Now, by definition of and Theorem 2.6, we have that is IT(0) for every . In particular, for any integer we have the following vanishings:
For we automatically have and hence is -regular with respect to as soon as
for which it is enough to have In particular, we obtain that
@@
Passing to the limit we get
and therefore
@@ 5. Gauss-Wahl maps
In particular, the right arrow is surjective if and only if separates -jets at every point of In this way we also see that and is the kernel of the evaluation map (which in [cohrank] is denoted by ).
@@
where the surjectivity at the bottom follows from the fact that the sheaf is supported on the diagonal (for example, see [Ricolfi, Proposition 1.3] for details). In particular, from the 5-lemma it follows that is surjective as soon as the evaluation map is surjective, that is, as soon as separates -jets at every point of
@@
Let be an ample line bundle on an abelian variety of dimension . Then:
@@
If separates -jets at every point then we have for every and while
for with equality if
@@
a) Write with and Without loss of generality we may assume that is symmetric and hence Applying to the exact sequences
@@
b) Now, write with a positive integer and . As separates -jets at every point, the sequence in (5.1) is also right exact. As for applying and taking cohomology we obtain then the following isomorphisms
where the last equality comes from part a). Moreover, if then and hence the above isomorphism also holds for as soon as Now, if then and therefore we have the following exact sequence:
where the group at the right is zero if The result then follows from part a).
@@
Note that, when separates -jets at every point, the above lemma tells us that is the vanishing threshold (see Definition 3.1) and, in this context, Note also that the invariant is already considered in [cohrank], where it is denoted by In the cited reference it is shown that there is a relation between such number and the base-point freeness threshold, when is globally generated. In the following we compute to show that there is an analagous relation between and
@@
and if and are algebraically equivalent then
@@
By the transformation formula (Proposition 2.11) it follows that for we have
@@
From Remark 3.10 we see that if is an ample line bundle then separates -jets at every point. Therefore, by Theorem 5.5 we have
@@
In particular the limit exists as an element of and by Proposition 4.1 it follows that
@@
There exists a principally polarized elliptic curve and a -dimensional polarized abelian variety such that
@@
The equivalence between a) and b) follows as a combination of Theorem 3.7 and Theorem 4.1. From Corollary 5.8 above, the equivalence a)c) is immediate. Finally, the equivalence a)d) is the content of Nakamaye’s theorem [Nakamaye, Theorem 1.1].